OpenStudy (jskinny0831):

What is the value of y so that AB with endpoints A(1, y) and B(6, −1) is parallel to CD with points C(1,2) and D(-4,-2)??

OpenStudy (jskinny0831):

can someone help me? i just need the equation on how to do it. I've been stuck on this for an hour

OpenStudy (faiqraees):

Okay whenever the direction between two pair of points is parallel. The gradient of both of them is same

OpenStudy (anonymous):

Parallel lines have the same slope. The slope equation is \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\] Plug int he values for each of your segments and set the slopes equal to solve for your "y"

OpenStudy (faiqraees):

So \[\large\rm \frac{y_2-y_1}{x_2-x_1} =\frac{y_3-y_4}{x_3-x_4} \]

OpenStudy (jskinny0831):

okay so -1-y/6-1 = -2-2/-4-1 ?

OpenStudy (anonymous):

yes

OpenStudy (faiqraees):

yep

OpenStudy (jskinny0831):

so how do i get y? im still kinda confused

OpenStudy (faiqraees):

Solve it

OpenStudy (faiqraees):

-1-y/5 = -4/-5

OpenStudy (jskinny0831):

-1-y/5=-4/-5

OpenStudy (jskinny0831):

yeah i got that

OpenStudy (anonymous):

then cross multiply

OpenStudy (anonymous):

-5(-1 - y) = -4(5)

OpenStudy (jskinny0831):

-5-5y=-20

OpenStudy (jskinny0831):

-5y=-20+5

OpenStudy (jskinny0831):

-5y=-15

OpenStudy (anonymous):

whoa hold up

OpenStudy (jskinny0831):

oh wait am i doing it wrong..?

OpenStudy (anonymous):

you just made a mistake distributing. it should be 5 - 5y = -20

OpenStudy (jskinny0831):

ohh i see where i made the mistake

OpenStudy (jskinny0831):

so -5y=-20-5

OpenStudy (jskinny0831):

-5y=-25

OpenStudy (jskinny0831):

-25/-5

OpenStudy (jskinny0831):

y=5?

OpenStudy (anonymous):

yep

OpenStudy (jskinny0831):

thank you so much!

OpenStudy (faiqraees):

(-1-y/5) x 5=(-4/-5)x5 -1-y = -4/-1 -1-y = 4 -y = 5 y = -5

OpenStudy (jskinny0831):

wait so i did it wrong?

OpenStudy (faiqraees):

You did a mistake -5(-1 - y) = -4(5) 5+5y = -20

OpenStudy (anonymous):

oh @FaiqRaees is right. I messed up the distribution as well

OpenStudy (jskinny0831):

ohhh lol ok thank you so much!!!